# Stellar reddening map from DESI imaging and spectroscopy

### v1

@@ 4.2 Validating the maps using galaxies as back light
The observations of galaxies are also affected by the dust of our galaxy. While it is more difficult to determine the intrinsic spectra of galaxies and estimate the galaxy reddening directly (although possible, e.g., see Peek & Graves 2010), we can exploit the fact that the distribution of galaxies is isotropic at large scales, and use galaxies as a uniform back light to test the reddening maps. Here we use the DESI emission line galaxy (ELG) sample (Raichoor et al., 2023) as the back light sample to test the DESI and SFD maps.
@@ 4.3 Validating the extinction curve with reddening color-color diagram
We also check for spatial deviations from the reddening curve by comparing the reddening in the two colors, $g-r$ and $r-z$. However, this difference is degenerate with zero point offsets in the $grz$ bands due to photometric calibration errors in the LS DR9 imaging data, and we find that the measured difference can be mostly explained with the zero point offsets measured with Gaia. We discuss this result in Section 5.1. The zero point uncertainties make it difficult to interpret the residual differences or make quantitative assessments. Nevertheless, we do not see deviations from the universal extinction curve that are significant compared to the photometric calibration errors within our survey footprint.
@@ 5.1 Photometric zero point uncertainties
Figure 15: Top panel: the $r-z$ reddening map. Middle panel: the difference between the $E(B-V)$ map from $g-r$ and the $E(B-V)$ map from $r-z$. Bottom panel: the difference map after removing the zero point offsets measured from Gaia spectrophotometry. Most of the features in the difference map are due to zero point systematics as they disappear after the Gaia correction. The stripes in the “corrected” difference map are due to systematics in the Gaia spectrophotometry. All three maps have NSIDE=128.
@@ 5.3 Dependence of the synthetic colors on the amount of extinction
For the reddening measurements, we have assumed that the synthetic colors from the best-fit rvspecfit models do not depend on the actual amount of extinction of the observed stars. To verify this assumption, we perform the following test. We randomly select a total of 99 stars that sample the different observing programs and target classes, and artificially add (or remove) extinction by multiplying (or dividing) the observed spectra with the additional dust attenuation (assuming the Fitzpatrick 1999 extinction curve). For each star, we generate new “observed” spectra with “added”/“removed” extinction $\delta E(B-V)$ in a 0.05 mag-spaced grid while restricting the “new” $E(B-V)$ (which is $E(B-V)_{\mathrm{SFD}}+\delta E(B-V)$) to within -0.05 to 0.25. We also modify the inverse variance of the fluxes such that the S/N is unchanged. We run rvspecfit on the artificially reddened/dereddened spectra, and compare the new synthetic $g-r$ color with the original synthetic color.
@@ 6 Data availability
The observed spectra, rvspecfit synthetic spectra, and catalogs containing the stellar parameters and synthetic photometry will be released along with DESI Y1 Data Release (which will include the first-year data) and Y3 Data Release (which will include the remainder of the data, i.e., data from the “daily” reduction, used in this paper). The synthetic spectra that will be made public can be used to measure the reddening using imaging data from a different survey (potentially with better photometric calibration).
@@ 7 Conclusion
We would like to thank Douglas Finkbeiner for helpful discussions during preparations of the dust map. R.Z. is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE–AC02–05CH11231. This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
@@
Software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), HEALPix/HEALpy (Górski et al., 2005; Zonca et al., 2019), Matplotlib (Hunter, 2007), Numpy (Harris et al., 2020), statsmodels (Seabold & Perktold, 2010)
@@ Appendix A Gaia-to-LS transformation
We use stars with $1816<G<$ 18 for the fit, and we apply the imaging quality cuts similar to those in Table 2. The fitting is done separately for the Northern and Southern imaging. We use 12-th order polynomials for the $g$ and $r$ bands and 7-th order for the $z$ band, using the Robust Linear Models (RLM) routine from statsmodels. The scatter, in normalized median absolute deviation, in the predicted LS magnitudes is 18, 11, 17 mmag in $g$, $r$, $z$ bands, respectively (the numbers are similar for the DESI stars used in the reddening measurements). The Gaia-predicted LS magnitudes have been used in the bright star “halo” subtraction in the Legacy Surveys pipeline since DR9777https://www.legacysurvey.org/dr9/psf/, and the high polynomials are necessary to accurately predict the LS magnitudes of the full range of stars (without color cuts). The DESI stars used for reddening only occupy a narrow range in $G_{\mathrm{BP}}-G_{\mathrm{RP}}$ color. Figure 17 shows the transformations to the DECam filters.

### v3

@@ 4.2 Validating the maps using galaxies as back light
The observations of galaxies are also affected by the dust of our galaxy. While it is more difficult to determine the intrinsic spectra of galaxies and estimate the galaxy reddening directly (although possible, e.g., see Peek & Graves 2010), we can exploit the fact that the distribution of galaxies is isotropic at large scales, and use galaxies as a uniform back light to test the reddening maps. It is worth noting that the pioneering work of Burstein & Heiles (1982) made use of galaxy counts to produce a reddening map that was widely used before the SFD map became available. Here we use the DESI emission line galaxy (ELG) sample (Raichoor et al., 2023) as the back light sample to test the DESI and SFD maps.
@@ 4.3 Validating the extinction curve with reddening color-color diagram
Previous studies, such as Schlafly et al. (2016) and Zhang & Green (2025), have measured significant spatial variations in $R_{\mathrm{V}}$, but they are limited to high-extinction regions at low Galactic latitudes where $R_{\mathrm{V}}$ can be more easily measured. Here we check for spatial deviations in the reddening curve by comparing the reddening in the two colors, $g-r$ and $r-z$; see Figure 15. However, this difference is degenerate with zero point offsets in the $grz$ bands due to photometric calibration errors in the LS DR9 imaging data, and we find that the measured difference can be mostly explained with the zero point offsets measured with Gaia. We discuss the zero point offsets in Section 5.1. The zero point uncertainties make it difficult to interpret the residual differences or make quantitative assessments. Nevertheless, we do not see deviations from the universal extinction curve that are significant compared to the photometric calibration errors within our survey footprint.
@@ 5.1 Photometric zero point uncertainties
Figure 15: Top panel: the $r-z$ reddening map. Middle panel: the difference between the $E(B-V)$ map from $g-r$ and the $E(B-V)$ map from $r-z$. Bottom panel: the difference map after removing the zero point offsets measured from Gaia spectrophotometry. Most of the features in the difference map are due to zero point systematics as they disappear after the Gaia correction. The stripes in the “corrected” difference map are due to systematics in the Gaia spectrophotometry. All three maps have NSIDE=128. Note that the bottom panel is only shown for illustration, and we do not apply the Gaia correction to our maps and catalogs.
@@ 5.3 Dependence of the synthetic colors on the amount of extinction
For the reddening measurements, we have assumed that the synthetic colors from the best-fit rvspecfit models do not depend on the actual amount of extinction of the observed stars. To verify this assumption, we perform the following test. We randomly select a total of 99 stars that sample the different observing programs and target classes, and artificially add (or remove) extinction by multiplying (or dividing) the observed spectra with the additional dust extinction (assuming the Fitzpatrick 1999 extinction curve). For each star, we generate new “observed” spectra with “added”/“removed” extinction $\delta E(B-V)$ in a 0.05 mag-spaced grid while restricting the “new” $E(B-V)$ (which is $E(B-V)_{\mathrm{SFD}}+\delta E(B-V)$) to within -0.05 to 0.25. We also modify the inverse variance of the fluxes such that the S/N is unchanged. We run rvspecfit on the artificially reddened/dereddened spectra, and compare the new synthetic $g-r$ color with the original synthetic color.
@@ 6 Data availability
The first-year portion of the observed spectra, rvspecfit synthetic spectra, and catalogs containing the stellar parameters and synthetic photometry have been released as Value-added Catalogs in DESI Data Release 1 (DESI Collaboration et al., 2025) [^9] https://data.desi.lbl.gov/doc/releases/dr1/vac/stellar-reddening/. The synthetic spectra can be used to measure the reddening using imaging data from other surveys and in different filters (potentially with better photometric calibration). The remainder of the data (taken during the second year of observations and processed in the “daily” reduction) will be made public with DESI Data Release 2.
@@ 7 Conclusion
We would like to thank Douglas Finkbeiner for helpful discussions during preparations of the dust map, and Leandro Beraldo e Silva, Carlos Allende Prieto and the anonymous referee for their helpful comments on the draft. R.Z. is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE–AC02–05CH11231. This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies.
@@
Software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), HEALPix/HEALpy (Górski et al., 2005; Zonca et al., 2019), Matplotlib (Hunter, 2007), Numpy (Harris et al., 2020), statsmodels (Seabold & Perktold, 2010), speclite (Kirkby et al., 2024), redrock (Bailey et al., 2025), rvspecfit (Koposov, 2019; Koposov et al., 2011).
@@ Appendix A Gaia-to-LS transformation
We use stars with $1816<G<$ 18 for the fit, and we apply the imaging quality cuts similar to those in Table 2. The fitting is done separately for the Northern and Southern imaging. We use 12-th order polynomials for the $g$ and $r$ bands and 7-th order for the $z$ band, using the Robust Linear Models (RLM) routine from statsmodels. The scatter, in normalized median absolute deviation, in the predicted LS magnitudes is 18, 11, 17 mmag in $g$, $r$, $z$ bands, respectively (the numbers are similar for the DESI stars used in the reddening measurements). The Gaia-predicted LS magnitudes have been used in the bright star “halo” subtraction in the Legacy Surveys pipeline since DR9101010https://www.legacysurvey.org/dr9/psf/, and the high polynomials are necessary to accurately predict the LS magnitudes of the full range of stars (without color cuts). The DESI stars used for reddening only occupy a narrow range in $G_{\mathrm{BP}}-G_{\mathrm{RP}}$ color. Figure 17 shows the transformations to the DECam filters.